OpenStudy (vshiroky):

Find the equation of the line that passes through (5,3) and is perpendicular to y=5x-6

OpenStudy (vshiroky):

@johnweldon1993 is it y=5x-22

OpenStudy (johnweldon1993):

Not quite...since I see perpendicular that means your slope should be the negative reciprocal of the other lines slope...so it's slope should be -1/5x

OpenStudy (vshiroky):

Dangit.

OpenStudy (vshiroky):

I may do another assignment since I am already going to be late to work haha. Will you be on for the next hour incase I need help?

OpenStudy (johnweldon1993):

So with that slope and that point, use the point slope form \[\large y - y_0 = m(x - x_0)\] putting in your numbers we have \[\large y - 3 = -\frac{1}{5}(x - 5)\] \[\large y = -\frac{1}{5}x + 4\] Would be your equation

OpenStudy (johnweldon1993):

and no sorry I have to head to work in a few minutes!

OpenStudy (vshiroky):

Ah ok thank you.

OpenStudy (johnweldon1993):

Glad to help :) anytime :)